OpenStudy (anonymous):

what is the area in square milimeters of a right trangle with a hypotenuse of 68 miilmeters and a leg of 60 milimeters?

OpenStudy (jack1):

area = 1/2 base x ht

OpenStudy (jack1):

use pythagorus theorum to find height

OpenStudy (jack1):

u understand pythag theorum...?@linamireles16

OpenStudy (anonymous):

no still dont understand

OpenStudy (jack1):

ok, no worries you have a right angled triangle, so you're able to use the formula: \[h^2 = a^2 + b^2\] where h = hypotenuse a = ht b = base 

OpenStudy (jack1):

so use h = 68 and b = 60 in the above formula to find a

OpenStudy (jack1):

then use a nd b in: area = 1/2 base x ht to find answer

OpenStudy (jack1):

so how'd you go for a?

OpenStudy (anonymous):

idid it wrong

OpenStudy (anonymous):

i got 1024

OpenStudy (jack1):

ok... what did u do?

OpenStudy (anonymous):

for b

OpenStudy (jack1):

thats fine, a^2 = 1024, so what does a = ??

OpenStudy (anonymous):

4624=3600+bsquares

OpenStudy (anonymous):

sqare rooredd it and got 36 which is not a valid answer

OpenStudy (jack1):

nah, sqrt 1024 = 32 hey

OpenStudy (anonymous):

yea you know what i mean ah any how all my answers are 0000 options

OpenStudy (anonymous):

to the thousand

OpenStudy (jack1):

that's cool, you now have a and b, so find area area of triangle = 1/2 x a x b

OpenStudy (jack1):

a = 32 b = 60

OpenStudy (anonymous):

my aswer choices are 2,040 1,920 1,088 960

OpenStudy (jack1):

area of triangle = 1/2 x a x b a = 32 b = 60 u tell me which is right dude?

OpenStudy (jack1):

\[area = (a \times b) / 2\]

OpenStudy (anonymous):

so its 32x60/2?

OpenStudy (jack1):

yep

OpenStudy (anonymous):

thanks i got it

OpenStudy (anonymous):

what about 65 meters and leg of 25 meters?

OpenStudy (jack1):

sweet, welcome

OpenStudy (jack1):

as for 2nd q, do the same thing find the verticle height and base length use that to calculate area

OpenStudy (anonymous):

is it 750?

OpenStudy (jack1):

so if you have right angle triangle, one side = base length one side = verticle height

OpenStudy (jack1):

depends if 65 is hypotenuse or verticle height...?

OpenStudy (anonymous):

hypotenuse

OpenStudy (jack1):

cool, so what was vertical height length then...?

OpenStudy (anonymous):

idk i just guest

OpenStudy (jack1):

ok... might work but probly not the best approach hey

OpenStudy (anonymous):

i got it right

OpenStudy (jack1):

noice! good guess dude

OpenStudy (anonymous):

anway whats the approxitmate side lenght of a square with a diagram of 16